Let A = {(x, y, z) E R^3ㅣx - y + z = 0} where E represents "an element of" and R the set of real numbers
A mapping f is defined in R^(2*2) , the set of all square matrices of order 2 with entries in R by f: R^(2*2) → R^2
A → Av where v = (-1 2).
(a) Determine Im F. Give a basis of it and precise its dimension.
1
Expert's answer
2021-02-12T10:29:27-0500

Let "f: \mathbb R^{2*2} → \mathbb R^2", "A \u2192 Av" where "v = (-1, 2)" . Let us show that "Im f=\mathbb R^2", that is, "f" is surjective. Fix any "(a,b)\in\mathbb R^2". Then for the matrix "A=\left(\begin{array}{cc} -a & 0 \\ -b & 0\end{array}\right)" we have that "Av=\left(\begin{array}{cc} -a & 0 \\ -b & 0\end{array}\right) \left(\begin{array}{cc} -1 \\ 2\end{array}\right)=
\left(\begin{array}{cc} a \\ b\end{array}\right)", and therefore, "Im f=\mathbb R^2". The standard basis for "Im f=\mathbb R^2" consists of the linearly independent vectors "e_1=(1,0)" and "e_2=(0,1)". Thus, "\dim(Im f)=\dim(\mathbb R^2)=2."